### v2

Abstract
An article usually includes an abstract, a concise summary of the work covered at length in the main body of the article.
Usage
Secondary publications and information retrieval purposes.
Structure
[…]
Abstract
Over the last five years, McGill University’s Office of Science Education (OSE) has partnered with faculty members from the Department of Physics to form an education working group with the aim of charting the progression of students’ conceptual understanding of uncertainties across their undergraduate degree. The research conducted by this group seeks to provide further insight into both the experimental skillset that students gain through undergraduate laboratory courses and how the department could address noticeable gaps in student understanding. In this paper, we evaluate the conceptual understanding of uncertainty using the Concise Data Processing Assessment (CDPA) test. First, we characterize the physics laboratory curriculum at McGill University by evaluating the evolution of CDPA scores across consecutive laboratory courses, and further propose the utilization of this tool for identifying of gaps in student understanding. Following the analysis of student responses (N = 2023) we specifically investigate data collected in second-year courses to better understand what student errors reveals about common misconceptions in experimental physics. This more in-depth analysis focuses on data collected from students at the beginning and end of two consecutive experimental laboratory courses that build on each other. By the end of the second course, students have engaged with all the material covered in the CDPA test. Interestingly, we did not observe any change in CDPA scores throughout and post pandemic. We notice a marked upward shift in student understanding of uncertainties, however the results indicate that many students continue to struggle with uncertainties, basic data analysis and curve fitting.
I Introduction
Understanding measurements of physical quantities and uncertainties is an important skillset for evaluating and interpreting data in scientific research. In recent years, the consideration and discussion of uncertainties has increasingly become more present and relevant in physics undergraduate curricula [1, 2, 3], in line with the American Association of Physics Teachers (AAPT) recommendations for undergraduate physics labs [4]. Skills such as modeling [5], experimental design [6], data analysis, interpretation [7], and communication [8] are essential for the training of physicists. Developing a comprehensive understanding of how these competencies are scaffolded across an educational curricula can guide instructional changes and reinforce pedagogical practices that are further in line with the AAPT guidelines.
The study reported in this paper was conducted at McGill University, a research-intensive university in Quebec, Canada. Following an in-depth curriculum review process, the Chair of Physics Department invited a small group of professors teaching at different levels in the experimental physics program to partner with the University’s Office of Science Education (OSE) to form the physics education research (PER) group. Improving introductory, intermediate, and advanced undergraduate physics labs to reinforce the understanding of uncertainties was selected as a central point. This focus emerged through in-depth discussions about assessment conducted at peer institutions [9, 10, 11, 12, 13]. These studies highlight the importance of assessing the understanding of uncertainties as a means to revise and improve upon students’ experimental competencies in first-year laboratory courses [14, 15].
In the PER literature, a range of tests are applied to assess experimental competencies such as the Concise Data Processing Assessment (CDPA) [10], the Colorado Learning Attitudes about Science Survey for Experimental Physics (E-CLASS) [16], the Physics Measurement Questionnaire (PMQ) [17], the Laboratory Data Analysis Instrument (LDAI)[18], the Physics Lab Inventory of Critical Thinking (PLIC) [19], and more recently the Survey of Physics Reasoning on Uncertainty Concepts in Experiments (SPRUCE) [20]. Each of these evaluates different aspects of physics students’ learning. PLIC, for example, focuses on students’ evidence-based decision making by measuring abilities ranging from evaluation of data, methods, and conclusions to the proposal of next steps in hypothetical experiments. In comparison, SPRUCE and CDPA are examples of specific tools to assess the understanding of the concept of uncertainty in undergraduate labs. SPRUCE has a broader approach, identifying sources of uncertainty while an experiment is being conducted and leading up to the handling of uncertainty, whereas the CDPA is used specifically to assess error propagation skills, and the interpretation of equations, graphs, and uncertainties therein.
The CDPA was chosen by the McGill PER group as the primary investigative tool to conduct a targeted, longitudinal study across the physics department’s undergraduate laboratory courses. The objective was to chart the progression of students’ understanding of uncertainty over the course of their degree. In this paper, we present the results of a 4-year study in which we implemented the CDPA test across six courses, with a large dataset of N=2023 responses to the CDPA. Through this research, we focus on discussing relevant findings in fundamental PER topics that serve to guide the department’s current teaching goals and future pedagogical directions.
The objectives of this work are 1) to establish a validation of baseline CDPA data at McGill University compared to the data published in a previous study [10], 2) to investigate the progression of CDPA scores throughout different lab courses and across consecutive years, 3) to compare the evolution of student comprehension on the topic of uncertainty prior to, during, and directly following the global COVID-19 pandemic period (2019-2022) during which students experienced varying degrees of virtual versus in-person instruction, and finally 4) to propose a deeper interpretation of CDPA results in terms of mapping student abilities and identifying common student misconceptions. Where possible, we also provide additional contextual explanations of the features of our findings that may influence the data interpretation.
II Background
The CDPA was first implemented in 2011 at the University of British Columbia (UBC) and the University of Edinburgh [10]. It was initially developed for first-year physics major students in laboratory courses or in science-intensive programs that also required an experimental physics laboratory component. Graduate students and professors also participated in the study in order to ensure the validity of the test in measuring novice vs. expert performance and to further investigate its reliability when used by students at different levels of instruction. The CDPA test consists of 10 multiple choice questions, which vary in terms of difficulty, each with four possible answers. The questions cover a variety of learning outcomes such as identifying the adequacy of a fit to a given dataset, handling instrumental uncertainty, and inferring measurement units from the data. The test’s characteristics, such as its difficulty and its capacity to discriminate amongst the level of expertise of test-takers, are proposed and quantified in the original paper. Furthermore, the test is administered both as a pre-course and post-course assessment in order to reliably measure shifts in reasoning that can be linked to the learning outcomes of each respective course.
The results in [10] showed 1) that there are positive shifts in pre- to post- comparisons for the first-year courses, 2) that the test is sensitive to the level of expertise of the test-taker, and thus is able to distinguish between undergraduate students, graduate students, and professors, 3) that the test is suited to a broader public, a conclusion made based on the comparable results from UBC and the University of Edinburgh. All these findings are supported by statistical tests which confirm the reliability of the CDPA. For example, first-year UBC students scored an average of $3.91\pm 0.12$ (standard error of the mean, SEM) in the post-test, compared to an average of $5.31\pm 0.42$ (SEM) for the graduate students, and $8.00\pm 0.35$ (SEM) for professors. A similar, more marginal progression of knowledge was also shown across the undergraduate program by comparing the post-test of first-year students with that of fourth-year students who averaged $4.73\pm 0.26$ (SEM) in the assessment.
Since its publication, this ten-question test has stimulated discussions amongst the PER community with different proposed applications of the CDPA results. Namely, Day et al. [21] investigated the influence of gender in the assessment results, finding no evidence of gender differences in conceptual understanding of uncertainties. The CDPA was also the instrument used to quantify the role of guidance during the learning process and it was found that scaffolding student orientation improves performance compared to students that do not receive any guidance, and that there is a link between the invention process and performance [22]. Kontro, et al. [23] further explore the use of the CDPA for intermediate-level undergraduate physics courses and proposes the possibility of a relationship between CDPA final scores and E-CLASS results, though no significant correlation is observed.
III Methodology
III.1 Context
The findings presented in this paper span the academic years from 2019 to 2023. In response to the global pandemic in 2020, universities were forced to adopt an remote learning. Following this period, 2021 marked the beginning of a hybrid instructional format with classes held online, while lab sessions gradually transitioned back to an in-person format. In 2022, the laboratory classes covered in this study returned to a fully in-person format. For the purposes of this study, we consider that the contents of these courses did not change between the in-person to the online setting, since the courses followed the same learning outcomes in both cases. In keeping with the CDPA requirements, no adaptations were made to the original CDPA test and the entire ten-question test was provided to all student respondents. Furthermore, as requested by the authors, we did not reproduce or include direct citations of any of the ten CDPA questions in order to preserve the possible use of this assessment instrument for future research purposes. However, it must also be noted that, in order to ensure a higher proportion of student respondents, a bonus of $1\%$ in their final course grade was offered in each class if at least $80\%$ of students in a given course took the test.
Finally, to avoid statistically significant shifts due to students retaking the CDPA test, we made sure to ask student respondents to identify whether or not they had taken the CDPA test in a previous semester. Students with prior exposure to the CDPA test were omitted from this study. The test was conducted online through MS Forms and student respondents were instructed to use up to 30 minutes to complete it. The pre-tests were taken over the first three to four weeks of term, whereas post-tests were taken during the last three to four weeks of term. Tests without pre and post distincion were taken at the end of the term. We investigated a total of 8 courses: PHYS 101- Introduction to Mechanics, PHYS 131- Mechanics and Waves, PHYS 257- Experimental Methods I, PHYS 258- Experimental Methods II, PHYS 359- Honors Laboratory in Modern Physics I, PHYS 439- Majors Laboratory in Modern Physics, PHYS 469- Honors Laboratory in Modern Physics 2, and PHYS 434- Optics. All results presented here are summarized in Appendix appendix.
III.2 McGill Physics curriculum
The understanding of uncertainty in McGill Physics laboratory courses becomes progressively more important as the level of lab increases. This increasing importance is reflected in the progression of the weight assigned to students’ abilities to interpret and display their understanding of uncertainties in reports and graphs as they move closer to the end of their degree.
First-year students are often exposed to experimental physics through a laboratory portion of an introductory course if they are enrolled in a general science program or are pursuing a minor in physics. In these cases, at McGill University, students are required to take an introductory undergraduate laboratory course: PHYS 101- Introduction to Mechanics. The alternative for physics majors or honors students is to take part in the higher-level introductory course: PHYS 131- Mechanics and Waves. In these introductory courses, students explore basic mechanics concepts through simple confirmation experiments, where the concept of uncertainty is addressed but does not constitute a major portion of their final grade. Students who proceed to a physics-specific degree, i.e., majors and honors students, enroll during their second year in PHYS 257- Experimental Methods I and PHYS 258- Experimental Methods II, where they are formally introduced to uncertainties in experiments. In these two courses, which span one full academic year, students are required to report uncertainties and discuss the validity of their experimental results and findings based on their interpretation of these uncertainties.
By the conclusion of PHYS 257 and PHYS 258, students are expected to possess the skills needed to interpret, address, and visually represent measurements. Students who choose to focus on experimental physics take the 300- and 400-level courses, which are not part of the mandatory curriculum, and can therefore represent expertise-like level of understanding. The notation difference between 300 and 400 is simply a means to indicate which students are part of the honors program as opposed to the majors program, otherwise the curricular content covered in both PHYS 359 and PHYS 439 is identical. We did not consider a chart uniquely for PHYS 469, PHYS 434 and PHYS 439. We decided to join the PHYS 434, PHYS 439 and PHYS 469 courses to form the overall 400 level, because the sample size was too small and they correspond to similar levels of expertise.
III.3 Educational background
In the context of our university, situated in Québec, a notable portion of incoming freshman students exhibit diverse educational backgrounds. This diversity is attributed to the fact that some students have completed Collège d’Enseignement Général et Professionnel (CEGEP) before enrolling in university. Students who complete CEGEP and pursue a physics degree at univeristy skip the first year of the McGill curriculum, taking PHYS 257 and PHYS 258 as their initial university-level laboratory courses. During data collection and analysis, we could not discriminate whether students had completed CEGEP.
III.4 Physics Education Research (PER) Group at McGill
III.4.1 Origin of the PER group
The PER group who led for this study represented a partnership between the Department of Physics and the Office of Science Education (OSE). Following an in-depth curriculum review process, the chair of the department nominated five tenure track professors teaching at different levels in the experimental physics program to work with OSE staff (including the director, educational developers and a science education expert), the physics liaison librarian, as well as PhD students selected by the department. The goal was to measure the effectiveness of lab courses from a program perspective. Although membership evolved over the five-year period, the group met on a regular basis and operated similarly to a community of practice [cuddy2002cultivating] with distributed leadership and regular cycles of data collection and reflection.
III.4.2 PER professors feedback sessions
Following the most recent round of data collection, feedback sessions were held with the PER professors. The purpose of these sessions was to better understand the level of each of the department’s experimental courses and to map the specific abilities taught in each course as they related to the content of the CDPA questions. These sessions were used to situate each course in the McGill Physics undergraduate program based on their contents, and to better understand how students at various points in their degrees would be expected to respond to the CDPA. This analysis will be discussed further below.
III.5 Statistical analysis
Shapiro-Wilk [24] tests and Kolmogorov-Smirnov [25] test revealed non-normality in the distribution of CDPA scores, parametric tests were not suitable for between-group comparisons. Therefore, non-parametric tests were employed. Between-group differences in CDPA scores were assessed for statistical significance using the Kruskal-Wallis [26] tests. Pairwise comparisons were conducted to compare all groups and a $p$ value $<0.05$ was considered as statistically significant. For multiple comparisons, Dunn’s post hoc comparison test [27] was applied to adjust the $\alpha$-level as necessary. Bonferroni [28] correction was not applied. All statistical analyses were performed with GraphPad Prism ® (GraphPad Software Inc., San Diego, Calif., USA). All the graphs shown in this article display the mean as well as the standard error of the mean.
IV Results and Discussions
IV.1 Comparison to UBC
Establishing a comparison between universities across the country is useful to evaluate the progression of students at each level of their studies across academic years. In Figure 1, we present the comparisons between the results at McGill University (in blue) and the University of British Columbia (in blue), where the CDPA survey was first developed. At McGill, we consider the results from the year of 2019, comparing it to the reported results at UBC. The overall results across both universities showed an ability differentiate between novice and expert-level understanding as a function of a respondent’s level of education. We notice a maximum score for students in the fourth and final year of their undergraduate education, without significant change towards the graduate student level. As expected, professors from both universities score higher than all previous cohorts, but not a perfect score, which also points to to the high standard of difficulty of this test. These CDPA results allow us to draw an important conclusion: even at the higher end of the novice to expert scale, the average score is not a complete $10$, a feature consistent with Day et. al. [10].
Figure 1: Mean CDPA scores (left) for McGill University and (right) the University of British Columbia (UBC), as a function of level of program/degree stage in the year of 2019. The average score increases as a function of level of instruction.
IV.2 Longitudinal analysis of the McGill data
The total CDPA mean score for the courses investigated between 2019-2023 is presented on a scale from 0 to 10 in Figure 2a. From left to right, the order of level of expertise increases by course numbers. We notice statistically significant jumps in the CDPA scores when comparing the scores obtained during the very first mechanics course that contains a lab component, PHYS 101, to a more advanced laboratory, PHYS 257, when students are properly exposed to the concepts of uncertainty and the methods used to analyze them experimentally.
Moreover, in 2019 and 2020, we also analyzed pre- and post-test results for the initial first-year mechanics course, PHYS 131, to evaluate changes within the course because, throughout this course, students are asked to calculate uncertainties but are not specifically evaluated on that skill. In these two years, we notice that the CDPA scores obtained for the pre- and the post-test are similar, which is to be expected since students are only exposed to uncertainties through problem-solving. When comparing PHYS 131 with higher level courses, however, we do notice a statistically significant increase in the level of responses when compared to the experimental course PHYS 257 in 2019 and in 2020. In 2021, when only PHYS 131 post-test data was collected, we also notice a statistically significant jump between PHYS 131 and PHYS 258, which marks the increase in learning following undergraduate students’ second full year of experimental courses.
When comparing PHYS 257 to PHYS 258, we notice an increase in student understanding that is consistent across three out of four years under investigation. Most notably, however, the average CDPA score collected from students at the end of PHYS 359 is significantly superior to the results from all previous courses, whereas, for the 400-level courses, because the number of student respondents was smaller and their answers presenting a higher degree of deviation, we cannot make of the same claims.
In Figure 2b, we compare how the CDPA scores evolve for the same course as a function of time. Interestingly, for PHYS 101 and PHYS 131, the overall CDPA score has a combined average that does not change as a function of year, except in the rare case of PHYS 101 in 2022 where the score was abnormally low. It is important to note that no major changes were implemented in the course that could have explained this difference in score. In PHYS 257 and PHYS 258, the results show fluctuations that can be consistent with the learning curve faced at that level for each of the different cohorts. The structure, contents and evaluation methods used in these two courses did not change in any of the years under study. However, the difference is that in each year different instructors taught the course using the same syllabus . In the 2020 Winter session of PHYS 258, students only completed the laboratory component of the class which corresponds to 60 % of the syllabus, not accounting for the final project. That may partially explain the decrease in CDPA scores that year.
For higher level courses over the first three years, the overall score is also consistent with a jump observed in 2022. In that year, the number of students taking part in the class was smaller than in previous years but students still obtained significantly higher scores than previous cohorts. These findings are likely due to classroom dynamics that are notably different from the 200 level courses, which we will further argue plays an important role in students’ understanding of uncertainty.
Figure 2: (a) Bar graphs representing the total CDPA score out of 10 for first year (PHYS 101, PHYS 131), second year (PHYS 257 and PHYS 258), and third year and upper-level (PHYS 359, PHYS 400s, PHYS 434) undergraduate laboratory courses at McGill University, across four years (2019-2922). The stars (*) represent the degree of statistical significance between courses, and the errorbars refer to the standard error on the mean (SEM). (b) Total CDPA score out of 10 for the undergraduate laboratory courses, presented individually as a function of year.
The data presented in Figure 2 shows that for the two lower-level courses, PHYS 101 and PHYS 131, the overall CDPA score can be understood as being comparable to random guessing. This can be explained by the fact that, in those courses, students are introduced to the concept of uncertainties for the first time and are only capable of applying simple mechanics problems and experiments without a deeper conceptual understanding. The CDPA test evaluates student understanding in topics that involve uncertainty propagation, graph analysis with fits and fitting uncertainties, logarithmic fits, and evaluation of $\chi^{2}$. In the original CDPA paper [10], the authors discuss that the vast majority of students from introductory-level laboratory courses are not yet ready to answer all the survey questions in the first year of their undergraduate degree. This portrayal of the perceived accuracy of the CDPA measurement tool for introductory-level courses is in agreement with our observations and analysis of the understanding of uncertainty for students enrolled in PHYS 101 and PHYS 131.
In the following discussions, we hone in on the results obtained from the PHYS 257 and PHYS 258 courses because it is during this full academic year that students learn how to perform curve fitting, start considering uncertainties to assess the outcome of an experiment, and produce and interpret graphs based on measuring uncertainties. Comparisons at the higher level could also garner interesting results, however the population size of student respondents was not sufficiently large for each year of the study. Additionally, having pairwise comparisons between two courses within the same academic year allows for a better comparison in terms of the understanding of uncertainties. In the case of the McGill Physics curriculum, the two courses are complementary to one another and students’ abilities to understand concepts related to uncertainty is not developed until the end of the full academic year, as shown in Table 1. At the end of PHYS 258, students are asked to conduct a small research project in which they need to answer a research question of their choice. Students experimentally validate their measurements and study how the uncertainties should be considered in the analysis, as well as how to best mitigate errors.
In Table 1, we present the central concepts regarding uncertainties that are covered in each of the courses we investigate. While in PHYS 101 students know how to propagate uncertainties numerically and make graphs, it is not until PHYS 257 and PHYS 258 that they actually get to do more advanced experiments in which error analysis becomes crucial. In these two courses, the students’ experimental skills are put to the test and they are evaluated based on applications involving uncertainty propagation at different levels. In PHYS 257, students are exposed to and are required to assess systematic errors and perform error propagation. Moreover, in PHYS 258, they need to consider statistical uncertainties to get the best-fit curves, understand and justify their reasoning, and differentiate between different models. We present in Table 2 some example questions from the laboratory manuals at three different course levels.
Course Uncertainty propagation Experiments Analysis of uncertainties Uncertainty minimization Graph construction Graphs with uncertainties Grade for uncertainties Experimental design
PHYS 101 x x x x
PHYS 131 x x x x
PHYS 257 x x x x x x x
PHYS 258 x x x x x x x x
PHYS 359 x x x x x x x x
PHYS 400s x x x x x x x x
Table 1: Evaluation criteria related to uncertainties for each course in the study. Notably, since the first course, students are exposed to uncertainties and experiments. However, it is only at higher level laboratories with experiments that require a higher degree of precision that students need to evaluate the experimental uncertainties.
Course Example Questions
PHYS 101, PHYS 131 The value of the mass used in the experiment (measured with a scale) and its uncertainty. Discuss the uncertainty estimates. How did you measure uncertainties?
PHYS 257 Compare the standard deviation obtained from the manual measurements with the automated measurements made. Which measurement is more reproducible/precise? Comment on the difference. Assuming that systematic errors are absent, what effect does this have on the uncertainty in your measurement? Be quantitative in your answer, not vague and qualitative.
PHYS 258 For data analysis, we will convert our data tables into histograms. We will then fit Gaussian and Poissonian distributions to the histogram and compare $\chi^{2}$ and residual plots with each other.
Table 2: Example instructions taken from a laboratory manual of each lab. The treatment of uncertainties is mostly numerical, without a critical assessment. This changes in higher level labs, where students are asked to identify, interpret, and mitigate possible uncertainties faced experimentally.
IV.3 Pandemic results
The global COVID-19 pandemic of 2020 introduced a significant external factor that is separately considered in this work. As was earlier stated, although the syllabi of the courses included in this study did not change before, during, and after the pandemic, the formats in which these courses were delivered had to adapt due to provincial regulatory health guidelines. As such, even after smaller courses returned to an in-person teaching environment, many of the large-scale introductory courses, including PHYS 101, PHYS 131, PHYS 257, and PHYS 258 continued to be taught in either a hybrid or a fully remote format.
Surprisingly, however, we notice no significant difference in CDPA scores between these pre- and post-pandemic periods. As was pointed out by one of the course instructors, this lack of any statistically significant fluctuation in CDPA scores is, in and of itself, noteworthy. The consistency of these scores, as shown in Figures 2 and Figure 3, suggests that the pedagogical benefit of laboratory courses may have more to do with students gaining first-hand practical knowledge through manipulating laboratory instruments than with their conceptual understanding of measurement principles.
Furthermore, if exposure to experimental instruments is only part of the intended pedagogical value of laboratory courses, then how can these in-person labs be improved such that they enhance students’ understanding of concepts related to uncertainty to a degree that is distinguishable using a measurement instrument such as the CDPA? The consistency of these pandemic-era CDPA scores also raises the question whether the CDPA instrument can to distinguish between the impact of confirmation versus inquiry-guided instructional approaches. However, from the findings of this research, we deduce that in all undergraduate courses included in this study could benefit from more dedicated teaching of graphing and analysis of uncertainties, as we will discuss in the following sections.
Figure 3: CDPA average scores for PHYS 257 and PHYS 258, combined, in four periods: 1) pre-pandemic, 2) pandemic (with online classes), 3) pandemic hybrid (theoretical classes online, laboratory classes in -person), and 4) in person.
IV.4 Assessing student misconceptions
Wilson et al. [29] and Harrison et al. [30] describe the concept of uncertainty of physical measurements as a threshold concept in physics, emphasizing its critical role in deepening students’ understanding of various topics, including experimental data analysis and the fundamental principles of quantum physics [31]. Majidi’s review article further highlights that misconceptions about uncertainties can arise from three primary sources: the knowledge itself, teaching methods, or the students’ perspectives. Overall, misconceptions can be more detrimental than gaps in knowledge. This is because students tend to resist new information that conflicts with their existing beliefs, resulting in cognitive dissonance and frustration. Our ability to characterize misconceptions through the CDPA however is constrained to assessing the alternatives present in the test and mapping it to possible misconceptions.
Furthermore, Wilson et.al [29] discusses that understanding uncertainties can be divided into three different layers of understading: pattern recognition, formal procedure, and meaning. In this study, three skills were evaluated through interviews with students, and it was discussed that integrating the three elements can give a more complete view on student understanding. In terms of the CDPA, pattern recognition is applicable to questions that evaluate uncertainties in graphs, or numerical uncertainties. The second skill, formal procedure, is tested when students have to actually calculate the uncertainties for some of the numerical questions presented in the assessment. The skill related to the meaning of uncertainties cannot be directly assessed through the CDPA, but can potentially be inferred through the experimental outcomes of the laboratory activities.
By analyzing the results from novice to expert levels in two Canadian universities (Figure 1), we see that the CDPA is not an easy test. The distractors represented by each of the incorrect answers, which are linked to distinct numerical values and graphical representations, were not selected at random. Rather, they reflect a specific approach to reasoning that students could easily be led astray when answering the ten questions. Thus, in light of the characteristic challenges of this scheme, instead of interpreting the mean CDPA score as a final measurement, we propose a different approach to these results by focusing on the distribution of correct and incorrect responses for each question. We use this distribution as an additional layer of analysis that can be derived when using this instrument. The original CDPA paper divides the survey by specific abilities and classifies them in two distinct groupings. The first grouping, composed of Q1, Q2, Q5, Q6, Q9, and Q10 evaluates students’ understanding of the meaning and use of uncertainty in measurements. The second grouping, Q3, Q4, Q7 and Q8, measures their ability to relate functions, graphs, and numbers. In the analysis that follows, we chose to focus our attention on the first grouping in the analysis that follows since these questions specifically target the practical understanding of uncertainties. For further reference, a list of the ten questions, complete with their distractor answers and the justification for these answers, is available in the original paper [10].
In Figure 4, we present our results taken from PHYS 257 and PHYS 258 for two complete years of undergraduate experimental physics laboratory education. The pie charts show the correct answers highlighted in green and the incorrect answers represented in gray. Remarkably, we notice that the two different cohorts displayed in Figure 4, over two distinct years, show a very similar distribution of responses. This trend becomes particularly evident when narrowing in on the correct responses given by respondents from both PHYS 257 and PHYS 258. When comparing the wrong alternative answers one by one, we also identify similarities that might point to distinct approaches to student reasoning. Moreover, in comparing the progression from one course to another, we notice a positive increase towards the correct answer for most questions.
The professors we reported our findings to, each of whom had instructed the courses included in this study, were asked to analyze the descriptions and justifications of the detractor questions and concluded that they were in agreement with the parameters set by the original CDPA paper. After being presented with the range of answers gathered from their respective students, they were then prompted to evaluate the content of each incorrect answer to further examine the misconceptions associated with each possible alternative approach to reasoning. The instructors were then asked to classify whether students should be able to answer each question by the end of the course they instructed. Their responses were categorized as follows: 1) yes, these students should be able to answer the question, 2) no, these students should not be able to answer the question, or 3) maybe, a stronger student in this course might be able to answer the question but it is not a requirement to pass the course in question. The analysis for each question, indicated in Table 3 below, reflects the results of these feedback discussions. Overall, the mapping of misconceptions established by these McGill professors is in line with the previously published results.
Question Content
Should students
be able to answer
in PHYS 257?
Should students
be able to answer
in PHYS 258?
Possible misconception
Q1 Combining uncertainties of different measurements M Y Using weighted mean to evaluate different uncertainties
Q2 Expressing significant figures Y Y Displaying significant figures in text
Q5 Identifying the best-fit curve N M Prioritizing more points rather than the ones with smaller uncertainty on a fit
Q6 Identifying the best-fit curve N M Prioritizing more points rather than the ones with smaller uncertainty on a fit
Q9 Outcome of measurements and associated instrumental uncertainties M Y Evaluating different outcomes of measurements and their individual influence on a series
Q10 Combining uncertainties of different measurements M Y Uncertainty propagation when considering multiple experiments
Table 3: After discussion with the professors teaching in the labs/classes, this table shows which of the CDPA questions students from PHYS 257 and PHYS 258 should be able to answer (Y), might be able to answer (M) and should not be able to answer (N), with a description of the possible misconception linked to an incorrect response.
Figure 4: CDPA results for questions 1, 2, 5, 6, 9, and 10 for different years, and evolution as a function of level, comparing the post-test of PHYS 257 and PHYS 258. The correct multiple-choice answer is presented in green. The signs in Q2 and Q9 of PHYS 258 indicate no students chose alternatives d) in Q2 2020, and alternative e) in Q9 2021.
Figure 5: CDPA fraction of correct responses for each of the 10 questions for PHYS 359 in 2022.
We make a question by question analysis of the results, presented in Figure 5. When considering Q1, for example, students are still being exposed to the content and do not necessarily need to know how to answer the CDPA questions after completing PHYS 257. The only exception to this was possibly for a student with a stronger background of knowledge. The professors also explained that the incorrect alternatives show that students disregard uncertainties when taking an average value. There was further consensus that, by the end of PHYS 258, students should be able to identify and treat different uncertainties in measurements correctly.
Q2 presents the highest correct response rate of all the CDPA questions for this student demographic before the start of PHYS 257. This question addresses ability to understand significant figures in astronomical units. According to the professors, students should have entered the course already able to respond to this question. Incorrect alternatives do not present any specific insight into student reasoning other than a misconception on the expression of significant figures.
In contrast, the professors pointed out that only students who are very strong should be able to answer Q5 and Q6, and that their understanding of the concepts in these questions are presented and explored throughout PHYS 258. However, students are still not ready by the end of the course to compare different models for best fit, an approach they are only exposed to more in depth at a higher level. It must be reiterated that all the CDPA distractor answers are very close to the correct response and that, for these two questions, choosing an incorrect answer can be associated with disregarding numerical uncertainties graphically when considering a best fit line. For example, students often choose to fit the most points rather than fit the ones with the smallest uncertainty, an approach to reasoning that is indicative of their lack of knowledge on how to prioritize.
Most students answer Q9 correctly and they should be able to respond to it only by the end of PHYS 258 because, during PHYS 257, they are still being exposed to the content. In this case, the alternatives that are incorrect reflect a misunderstanding of statistics and how statistics impact experimental measurements.
Q10 investigates students’ capacity to combine numerical uncertainties from different measurements. The professors explain that they expect students to master this critical ability by the end of PHYS 258, as the students develop a better understanding of how to handle uncertainties through the progression of experiments they conduct during the course, culminating in a final project they complete.
IV.5 What the CDPA reveals about upper level lab courses
Having established a clearer picture of what the CDPA reveals about students’ understanding and misunderstanding of concepts related to uncertainty during their first comprehensive year of introductory lab courses, we now turn to what the CDPA can tell us about the level of learning that takes place in more advanced laboratory courses. As previously explained, beyond the 200-level courses, this study also includes data collected at the 300- and 400-levels, most notably PHYS 359 - Advanced Physics Laboratory 1, reserved for Honors Physics undergraduate students, and PHYS 439 - Majors Laboratory in Modern Physics, the equivalent course for Physics majors.
Between 2019 and 2022, the high variance in CDPA scores for PHYS 258 demonstrates that, at this point in their learning, students still do not have a complete mastery of the skills needed to evaluate experimental uncertainties. For PHYS 359, however, we note a significant increase in the CDPA score in 2022 that warrants a more in-depth investigation into what factors may have contributed to this leap in student understanding. As can be seen in Figure 2, the 2022 mean CDPA score was significantly higher, from a class average of $2.5\pm 2.0$ by the end of PHYS 258 to $4.5\pm 1.0$ by the end of PHYS 359. Although we can expect noticeable improvement in the CDPA score after this upper-level laboratory course, the extent of this statistically significant increase over one semester left us with more questions.
Through discussion with the instructor who taught PHYS 359 during the Winter 2022 semester, we could ascertain that certain structural elements of the course contributed to this demonstrable increase in student understanding of uncertainties. Three key factors stand out in the methods of the instruction and assessment of that semester. In terms of how the course was taught, the most significant difference between PHYS 258 and PHYS 359 is the class size. While PHYS 258 had over 100 students, the smaller enrollment numbers in PHYS 359 allowed for greater one on one interaction between students and their instructor. Furthermore, students also spend more time in the lab per week, and perform experiments that require a higher degree of precision and attention to uncertainties.
The course has been going through changes and adaptations. One example is that a course instructor for PHYS 359 in 2022 identified a noticeable weakness in incoming students’ level of understanding of uncertainties. This led to the development and implementation of an additional module dedicated to curve fitting in a self-paced Python tutorial. This module offered students more advanced exercises emphasizing graph interpretation and analysis, incorporating an interactive simulation feature and enabling personalized feedback. The teaching tool allowed the instructor to leverage the smaller class size for a more individualized instructional approach, dedicating time to helping students apply concepts related to uncertainties. With this personalized tool, the instructor could identify and correct student misconceptions more precisely.
Lastly, in addition to this novel instructional approach, the assessment scheme of PHYS 359 allocated 20 percent of the lab report grade, to skills such as justified quantitative statistical uncertainties, $\chi^{2}$ and residuals discussed quantitatively. As a larger portion of the course grade depended on students demonstrating their mastery of these concepts, their understanding of uncertainties in applied contexts was further strengthened.
V Conclusions
The findings presented in this study, which stemmed from a five year partnership between McGill University’s Office of Science Education (OSE) and professors from the Physics Department, represent the longest running implementation of the CDPA test. As such, this research provides particular insight into how longitudinal research using the CDPA instrument can serve to support pedagogical research in individual laboratory courses and across departmental curricula. To our knowledge, this is the longest-running CDPA survey, and the largest dataset, with a total of N=2023 students.
By comparing average CDPA scores from pre and post tests across all core laboratory courses in the McGill experimental physics undergraduate curriculum, we were able to distinguish distinct patterns in students’ understanding of concepts related to uncertainties from introductory 100-level courses through to upper-level 300-level and 400-level courses. As a general observation, our results show that, at the 100-level, students’ CDPA scores remained quite low and were mostly on-par with random guesswork. These results confirm UBC’s findings for introductory-level courses, which can be explained by the fact that, in first-year courses, students are introduced to the concept of uncertainties for the first time and are only expected to solve simple mechanics problems and conduct experiments without any deeper conceptual understanding.
At the 200-level, however, we notice an increase in students’ CDPA scores that remained largely consistent across all four years of this study. This uptick in student understanding can be explained by the fact that it is during this full academic year of laboratory courses that students learn how to perform curve fitting, start considering uncertainties to assess the outcome of an experiment, and produce and interpret graphs based on the measurement of uncertainties. Whereas in PHYS 257, students are asked to assess systematic errors and perform error propagation, in PHYS 258, they need to consider statistical uncertainties to get the best-fit curves, understand and justify their reasoning, and differentiate between different models. There is additional pedagogical value in the fact that PHYS 258 culminates in a small research project in which students are required to answer a research question of their choice, experimentally validate their measurements and demonstrate understanding of how the uncertainties should be considered in the analysis. Nevertheless, it is not until the 300-level and 400-level courses, when 20 percent of the lab report grade is allocated to justifying quantitative statistical uncertainties, that students’ demonstrated understanding of uncertainties develops into holistic concepts. In addition to using the CDPA instrument as a means to rank student understanding across the departmental laboratory curricula, our research also shows that this tool can be used to help instructors comprehend the nature of their students’ misconceptions. As shown in the breakdown of CDPA incorrect responses, Figure 4, CDPA results can be used as a remediation tool to help professors identify gaps in student understanding, thereby providing an additional means to refine their teaching approach for subsequent cohorts.
In conclusion, through this collaborative study, we were able to highlight new opportunities for pedagogical and structural improvement. We show how a better understanding of students misconceptions can be used enhances the professors’ understanding on student learning and improvements in a course-by-course basis. This research reinforces the value of collecting program level data and reporting the findings back to course professors for discussion. Furthermore, without engaging the professors in the process of data analysis, we would not have been able to pinpoint an educational benefit featured in this research, which can be tied to a simulation exercise developed by one professor that allows students to individually probe graphing with uncertainties. This type of resource, which could be adapted to pinpoint the learning needs of students across different levels of laboratory courses, shows promise to impact students’ conceptual understanding of uncertainties in addition to the existing course curriculum. Even without any additional resources, this study also points to a potential need for more inquiry-guided learning in laboratory courses, or, at the very least, it serves as an impetus to carve out more personalized attention in large-scale introductory laboratory courses.
appendix
Table of collected data throughout the years
All the collected data in this study is organized in Table 4, presented below.
Class Term Format Total Tested
PHYS 101- Introduction to Mechanics Fall 2020 PRE Online 475 452
Fall 2021 PRE Online 77 34
Fall 2022 In person 82 22
Total Tested (PHYS 101) N = 508
PHYS 131- Mechanics and Waves Fall 2019 PRE In person 177 167
Fall 2019 POST In person 323 313
Fall 2020 PRE Online 335 320
Fall 2020 POST Online 191 186
Fall 2021 Online 78 37
Fall 2020 Online 52 32
Total Tested (PHYS 131) N = 1055
PHYS 257- Experimental Methods I Fall 2019 In person 105 104
Fall 2020 Online 74 71
Fall 2021 In person and online 65 64
Fall 2022 In person 50 43
Total Tested (PHYS 257) N = 282
PHYS 258- Experimental Methods II Winter 2020 In person and online 44 44
Winter 2021 Online 39 38
Winter 2022 In person and online 44 40
Winter 2023 In person 23 19
Total Tested (PHYS 258) N = 141
PHYS 359 - Honors Laboratory in Modern Physics I 2023 In person 23 19
Total Tested (PHYS 359) N = 19
PHYS 434- Optics Fall 2020 Online 17 16
Total Tested (PHYS 434) N = 16
PHYS 469-Honors Laboratory in Modern Physics 2 Fall 2020 Online 2 2
Total Tested (PHYS 469) N = 2
Table 4: Summary of class formats, and testing for the various physics courses tested using the CDPA across different terms. The total refers to the number of students who answered the survey in that specific year, willing to participate or not in the survey.